OpenStudy (anonymous):

Determine the value of X if (211) base x = (152) base 8, how to do this?

OpenStudy (summiya):

X=7

OpenStudy (summiya):

1. convert 152 into decimal, we get 1*8+5*8+2*8=(106) base 10 2. convert 106 into base 7 & we get (211) base 7

OpenStudy (summiya):

(211)base x=(152)base 8 ..... eq. (1) convert into decimal both sides, we get 2x^2+x+1=106 this is quadratic equation,so 2x^2+x-105=0 and we get x1=7, x2=-7.5 consider x=7 and u get both sides equal http://www.math.com/students/calculators/source/quadratic.htm use to find value of quadratic equation online (:

OpenStudy (anonymous):

thank